OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−7, 5) and a directrix of y = −11

OpenStudy (anonymous):

what is half way between \((-7,5)\) and \(y=-11\)? that is the vertex and we need that first

OpenStudy (anonymous):



OpenStudy (anonymous):

should be clear that it is the point \((-7,-3)\) since \(\frac{5-11}{2}=-3\)

OpenStudy (anonymous):



OpenStudy (anonymous):

then use \[(x-h)^2=4p(y-k)\] with \[h=-7,k=-3\] and \(p\) is the distance between the vertex and the focus

OpenStudy (anonymous):

also hope it is clear that that distance is 8 so plug in the numbers and you are done we can check it if you like

OpenStudy (anonymous):

yw here is the check http://www.wolframalpha.com/input/?i=parabola+%28x%2B7%29^2%3D32%28y%2B3%29

hero (hero):

For any given focus \((x_1, y_1\)) and directrix \((x_2, y_2)\), such points can be inserted in to the formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) And simplified to produce the equation of the parabola. In this case you are given the focus\((-7, 5)\) and directrix \((x, -11)\). Inserting those points into the formula, you get: \((x - (-7))^2 + (y - 5)^2 = (x - x)^2 + (y - (-11))^2\) Which simplifies to \((x + 7)^2 + (y - 5)^2 = (y + 11)^2\) And expands to \(x^2 + 14x + 49 + y^2 - 10y + 25 = y^2 + 21y + 121\) Notice that \(y^2\) cancels on both sides which reduces the equation to \(x^2 + 14x + 49 -10y + 25 = 21y + 121\) \(x^2 + 14x + 49 + 25 - 121 = 21y + 10y\) \((x + 7)^2 - 96 = 31y\) Divide both sides by 31 to isolate \(y\)

hero (hero):

I made a mistake.

hero (hero):

You expand to get \(x^2 + 14x + 49 + y^2 - 10y + 25 = y^2 + 22y + 121\) \(y^2\) cancels to get: \(x^2 + 14x + 49 -10y + 25 = 22y + 121\) Which simplifes to \((x + 7)^2 - 96 = 32y\) Divide both sides by 32 to get \(\dfrac{(x + 7)^2}{32} - 3 = y\)

OpenStudy (anonymous):

wow !!

hero (hero):

It's easier than it looks @satellite73

hero (hero):

I'm not as fast as you with the Latex

OpenStudy (anonymous):

where in the world does this come from \[(x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\]?

OpenStudy (anonymous):

looks like the distance formula ?

hero (hero):

Yes

OpenStudy (anonymous):

i see but you don't need a distance formula when the two points are on the same vertical or horizontal axis i.e. i don't need the distance formula to find the distance between \((-7,5)\) and \((-7,-11)\) it is just 16 by your eyeballs

OpenStudy (anonymous):

not that it is wrong, it is just a lot of algebra for not much payoff imho

hero (hero):

I can do it much quicker by hand without the latex. It's not that much work. I was just showing the detailed steps.

hero (hero):

It's not that difficult to expand binomial squares.

hero (hero):

You prefer the graphical approach. I prefer the algebraic approach

OpenStudy (anonymous):

i agree, but i can't do it in my head i can find the distance between any two points on the same vertical or horizontal in my head though

hero (hero):

Most students prefer the algebraic approach though and avoid graphical methods at all costs.